A billiard ball is hit and travels in a straight line. If 𝑠 π‘π‘š is the distance of the ball from

its initial position at 𝑑 𝑠𝑒𝑐, then 𝑠 = 100𝑑2 + 100𝑑. If the ball hits the cushion that is 39 π‘π‘š from its

initial position, at what velocity does it hit the cushion?Β 


1
Expert's answer
2021-11-11T12:26:38-0500

Given:

"s(t) =100 t^{2}+100 t \\
s=39 \mathrm{~cm} . \\
\Rightarrow 39=100 t^{2}+100 t \\
\Rightarrow 100 t^{2}+100 t-39=0 \\
\Rightarrow \frac{-100 \pm \sqrt{(100)^{2}-4(-39)(100)}}{2(100)} \\
\Rightarrow \frac{-100 \pm \sqrt{10000+15600}}{200} \\
=\frac{-100 \pm 160}{200} \\
=\frac{-100+160}{200} \ [\text{t can not be negative}] \\ 
=\frac{60}{200}=\frac{3}{10}"

Velocity "(t)=\frac{d}{dt}[100t^2+100t]=200t+100\\"

At "t=\frac{3}{10}"

Velocity "=" "200(\frac{3}{10})+100=160\ cm/s"